# Fine-Grained Retrieval-Augmented Generation for Visual Question Answering

### v2

@@
Abstract
Visual Question Answering (VQA) focuses on providing answers to natural language questions by utilizing information from images. Although cutting-edge multimodal large language models (MLLMs) such as GPT-4o achieve strong performance on VQA tasks, they frequently fall short in accessing domain-specific or the latest knowledge. To mitigate this issue, retrieval-augmented generation (RAG) leveraging external knowledge bases (KBs), referred to as KB-VQA, emerges as a promising approach. Nevertheless, conventional unimodal retrieval techniques, which translate images into textual descriptions, often result in the loss of critical visual details. This study presents fine-grained knowledge units, which merge textual snippets with entity images stored in vector databases. Furthermore, we introduce a knowledge unit retrieval-augmented generation framework (KU-RAG) that integrates fine-grained retrieval with MLLMs. The proposed KU-RAG framework ensures precise retrieval of relevant knowledge and enhances reasoning capabilities through a knowledge correction chain. Experimental findings demonstrate that our approach significantly boosts the performance of leading KB-VQA methods, achieving an average improvement of approximately 3% and up to 11% in the best case.
Fine-Grained Retrieval-Augmented Generation for Visual Question Answering
 Zhengxuan Zhang1thanks:   Zhengxuan Zhang and Yin Wu contributed equally to this work., Yin Wu [^1] footnotemark: 1, Yuyu Luo1,2, Nan Tang1,2thanks:   Nan Tang is the corresponding author 1The Hong Kong University of Science and Technology (Guangzhou) 2The Hong Kong University of Science and Technology {zzhang393, ywu450}@connect.hkust-gz.edu.cn {yuyuluo, nantang}@hkust-gz.edu.cn
1 Introduction
Figure 1: Sample VQA Solution with MLLMs: (a) Direct answer without additional knowledge. (b) Single-modality coarse-grained RAG. (c) Our proposal KU-RAG.
@@ 1 Introduction
2 Preliminary
2.1 Task Definition
@@ 2.2 Knowledge Unit
$KU=\{K_{i},K_{n},K_{t}\}$ (1)
Figure 2: Illustration of general coarse-grained retrieval, fine-grained retrieval and knowledge unit retrieval for VQA.
@@
3 Knowledge Unit Management
3.1 Construction
Figure 3: The construction of knowledge units with entity and event. The entity list comes from the work of  Hu et al. (2023a), while the events are sourced from E-VQA (Yang et al., 2023).
Figure 4: Overview of KU-RAG Framework
@@ 3.1 Construction
$V_{c_{i}}=\text{Encoder}(c_{i})$ (2)
$D_{c}=(V_{c1},V_{c2},\dots,V_{cn})$ (3)
@@
$V_{i_{j}}=\text{Encoder}(i_{j})$ (4)
$D_{i}=(V_{i_{1}},V_{i_{2}},\dots,V_{i_{n}})$ (5)
3.2 Database-level Operation
@@
4 Knowledge Unit Retrieval Framework
@@ 4.1 Query Processing
$V_{q_{t}}=\text{Encoder}(Q_{t})$ (7)
$S_{n}=\text{Sim}(V_{q_{t}},V_{o_{i}})$ (8)
@@
$V_{q_{i}}=\text{Encoder}(Q_{i})$ (9)
4.2 Knowledge Unit Matching
@@
$KU=\text{Matching}(V_{q_{i}})$ (10)
$KU^{\prime}=\text{top-k}(KU)$ (11)
@@
$V_{q_{t}}^{\prime}=\text{Encode}(Q_{t}^{\prime})$ (13)
@@
$C^{\prime}=\text{top-k}(\text{Sim}(V_{q_{t}},C_{ku}))$ (14)
@@
5 MLLM Reasoning and Generation
@@
$MP=(I^{\prime}_{1},I^{\prime}_{2},\dots,I^{\prime}_{n})$ (15)
@@
$A=\text{MLLM}(MP,Prompt,(Q,A_{0}))$ (17)
@@ 6.1 Dataset
  • OVEN (Hu et al., 2023a): An Open-domain Visual Entity Recognition dataset, primarily examining the ability to recognize the names of visual entities.
@@
Table 1: Characteristics of Different Dataset
@@
Table 2: Main results of the experiment. And indicates that the result is from experiments conducted on the full version of the test set, sourced respectively from Hu et al. (2023a) and Chen et al. (2023). Except for the SOTAs, which are trained ([Uncaptioned image]), other methods are performed in a zero-shot scenario ([Uncaptioned image]).
@@ 6.2 Baseline
  • SOTA: For OVEN dataset, we use PaLI-17B (Chen et al., 2022), as reported by Hu et al. (2023a). For INFO SEEK and OK-VQA datasets, the SOTA model is PaLI-X (Chen et al., 2022), as reported in the work of Chen et al. (2023). For E-VQA dataset, we adopt the best results of the SOTA model MuKEA (Ding et al., 2022), as reported Yang et al. (2023).
@@
  • LlaVa NEXT-7b Liu et al. (2024): The 7b parameter version of the latest LlaVa model, an open-source MLLM.
@@ 6.4 Main Result
Figure 5: The Results of Ablation Study
6.5 Ablation Study
@@
7 Related Work
7.1 Knowledge-based Visual Question Answering
Knowledge-based Visual Question Answering (KB-VQA) aims to leverage external knowledge to assist in answering questions about images  (Marino et al., 2019; Gardères et al., 2020). In early KB-VQA approaches (Zhu et al., 2020; Gao et al., 2022; Lin and Byrne, 2022), Wikipedia was often used as the external knowledge source for KB-VQA Caffagni et al. (2024), leading to the common adoption of a retriever-reader framework. This framework first retrieves textual knowledge relevant to the question and image, and then “reads” the text to predict the answer. However, this passage dense retrieval method is a unimodal, coarse-grained text-to-text retrieval process, which struggles with specific, fine-grained questions such as visual entities (Hu et al., 2023a), events (Yang et al., 2023), and visual information-seeking questions (Chen et al., 2023).
Since the emergence of LLMs, some methods have explored using implicit knowledge from LLMs in addition to retrieving information from databases like Wikipedia Caffagni et al. (2024). Typically, they convert images into tags or captions and then use GPT to retrieve related knowledge (Gui et al., 2021). However, there is a gap between the query and the LLM’s knowledge source. To address this,  Hu et al. (2023b) proposed a prompt-guided image captioning method that controls the visual entities in generated captions based on textual queries, replacing general captions with question-dependent ones. Although some methods attempt to mitigate the loss of visual information by incorporating visual features (Salaberria et al., 2023) or enriching prompts with candidate answers (Shao et al., 2023), they have not completely solved the issue of information loss. Recently, Hao et al. (2024) introduce a self-bootstrapped visual-language model that refines retrieved knowledge using a selector-answerer framework, significantly improving knowledge selection and QA accuracy, but their method still requires complex training.
@@ 7.1 Knowledge-based Visual Question Answering
7.2 Multimodal Retrieval-augmented Generation

### v3

# Fine-Grained Knowledge Structuring and Retrieval for Visual Question Answering

@@
Abstract.
Visual Question Answering (VQA) focuses on providing answers to natural language questions by utilizing information from images. Although cutting-edge multimodal large language models (MLLMs) such as GPT-4o achieve strong performance on VQA tasks, they frequently fall short in accessing domain-specific or the latest knowledge. To mitigate this issue, retrieval-augmented generation (RAG) leveraging external knowledge bases (KBs), referred to as KB-VQA, emerges as a promising approach. Nevertheless, conventional unimodal retrieval techniques, which translate images into textual descriptions, often result in the loss of critical visual details. To address these challenges, this study presents two key innovations. First, we introduce fine-grained knowledge units that consist of multimodal data fragments (e.g., text fragments, entity images, and so on) in a structured manner. Rather than merely refining retrieval mechanisms, we prioritize the systematic organization and management of these knowledge units, ensuring that the structuring process itself enhances retrieval quality. Second, we propose a knowledge unit retrieval-augmented generation framework (KU-RAG) that seamlessly integrates fine-grained retrieval with MLLMs. Our KU-RAG framework not only ensures precise retrieval of relevant knowledge but also enhances reasoning capabilities through a knowledge correction chain. Experimental results demonstrate that our approach consistently outperforms existing KB-VQA methods across four benchmarks, achieving an average improvement of approximately 3% and up to 11% in the best case. The source code is available at: https://github.com/zzhang393/KU-RAG.
1. Introduction
Figure 1. Sample VQA Solution with MLLMs: (a) Direct answer without additional knowledge. (b) Single-modality coarse-grained RAG. (c) Our proposal KU-RAG.
@@ 1. Introduction
2. Preliminary
2.1. Task Definition
@@ 2.2. Knowledge Unit
(1) $KU=\{K_{i},K_{n},K_{t}\}$
@@
Figure 2. Illustration of general coarse-grained retrieval, fine-grained retrieval and knowledge unit retrieval for VQA.
3. Knowledge Unit Management
3.1. Construction
@@
Figure 3. The construction of knowledge units with entity and event. The entity list comes from the work of  Hu et al. (2023b), while the events are sourced from E-VQA (Yang et al., 2023).
@@
(2) $V_{c_{i}}=\text{Encoder}(c_{i})$
(3) $D_{c}=(V_{c1},V_{c2},\dots,V_{cn})$
@@
(4) $V_{i_{j}}=\text{Encoder}(i_{j})$
(5) $D_{i}=(V_{i_{1}},V_{i_{2}},\dots,V_{i_{n}})$
3.2. Database-level Operation
@@
4. Knowledge Unit Retrieval Framework
Figure 4. Overview of KU-RAG Framework
@@ 4.1. Query Processing
(7) $V_{q_{t}}=\text{Encoder}(Q_{t})$
(8) $S_{n}=\text{Sim}(V_{q_{t}},V_{o_{i}})$
@@
(9) $V_{q_{i}}=\text{Encoder}(Q_{i})$
4.2. Knowledge Unit Matching
@@
(10) $KU=\text{Matching}(V_{q_{i}})$
(11) $KU^{\prime}=\text{top-k}(KU)$
@@
(13) $V_{q_{t}}^{\prime}=\text{Encode}(Q_{t}^{\prime})$
@@
(14) $C^{\prime}=\text{top-k}(\text{Sim}(V_{q_{t}},C_{ku}))$
@@
5. MLLM Reasoning and Generation
@@
(15) $MP=(I^{\prime}_{1},I^{\prime}_{2},\dots,I^{\prime}_{n})$
@@
(17) $A=\text{MLLM}(MP,Prompt,(Q,A_{0}))$
@@ 6.1. Dataset
  • OVEN (Hu et al., 2023b): An Open-domain Visual Entity Recognition dataset, primarily examining the ability to recognize the names of visual entities.
@@
Table 1. Characteristics of Different Dataset
@@
Table 2. Main results of the experiment. And indicates that the result is from experiments conducted on the full version of the test set, sourced respectively from Hu et al. (2023b) and Chen et al. (2023). Except for the SOTAs, which are trained ([Uncaptioned image]), other methods are performed in a zero-shot scenario ([Uncaptioned image]).
@@ 6.2. Baseline
  • SOTA: For OVEN dataset, we use PaLI-17B (Chen et al., 2022), as reported by Hu et al. (2023b). For INFO SEEK and OK-VQA datasets, the SOTA model is PaLI-X (Chen et al., 2022), as reported in the work of Chen et al. (2023). For E-VQA dataset, we adopt the best results of the SOTA model MuKEA (Ding et al., 2022), as reported Yang et al. (2023).
@@
  • LlaVa NEXT-7b (Liu et al., 2024): The 7b parameter version of the latest LlaVa model, an open-source MLLM.
@@ 6.4. Main Result
Figure 5. The Results of Ablation Study
6.5. Ablation Study
@@
7. Related Work
7.1. Knowledge-based Visual Question Answering
Knowledge-based Visual Question Answering (KB-VQA) aims to leverage external knowledge to assist in answering questions about images  (Marino et al., 2019; Gardères et al., 2020). In early KB-VQA approaches (Zhu et al., 2020; Gao et al., 2022; Lin and Byrne, 2022), Wikipedia was often used as the external knowledge source for KB-VQA (Caffagni et al., 2024), leading to the common adoption of a retriever-reader framework. This framework first retrieves textual knowledge relevant to the question and image, and then “reads” the text to predict the answer. However, this passage dense retrieval method is a unimodal, coarse-grained text-to-text retrieval process, which struggles with specific, fine-grained questions such as visual entities (Hu et al., 2023b), events (Yang et al., 2023), and visual information-seeking questions (Chen et al., 2023).
Since the emergence of LLMs, some methods have explored using implicit knowledge from LLMs in addition to retrieving information from databases like Wikipedia (Caffagni et al., 2024). Typically, they convert images into tags or captions and then use GPT to retrieve related knowledge (Gui et al., 2021). However, there is a gap between the query and the LLM’s knowledge source. To address this,  Hu et al. (2023a) proposed a prompt-guided image captioning method that controls the visual entities in generated captions based on textual queries, replacing general captions with question-dependent ones. Although some methods attempt to mitigate the loss of visual information by incorporating visual features (Salaberria et al., 2023) or enriching prompts with candidate answers (Shao et al., 2023), they have not completely solved the issue of information loss. Recently, Hao et al. (2024) introduce a self-bootstrapped visual-language model that refines retrieved knowledge using a selector-answerer framework, significantly improving knowledge selection and QA accuracy, but their method still requires complex training.
@@ 7.1. Knowledge-based Visual Question Answering
7.2. Multimodal Retrieval-augmented Generation